# A Unified Framework for Forward and Inverse Problems in Subsurface Imaging using Latent Space Translations

### v1

@@ 5 Results and Discussions
Quantitative Comparisons: Figure 3(a) compares MAE on velocity predictions while solving the inverse problem across different OpenFWI datasets. We can see that our proposed models, which capture complex non-linear relationships in the latent space using U-Net and IU-Net architectures, consistently outperform baseline methods InversionNet, VelocityGAN, & Auto-Linear across majority of datasets. Overall, the large Latent U-Net model shows the best performance in inverse modeling followed by the small Latent U-Net model. On complex datasets such as Style family and CFB, Invertible X-Net with and without cycle loss are also comparable to the small Latent U-Net model. In most cases, the Invertible X-Net model trained with the cycle loss performs at par with the Invertible X-Net trained with supervised loss except in a few cases like CFA and FFA. In addition to MAE, we also evaluate baselines using other metrics reported in Table 7 in the Appendix.
Figure 3(b) compares the performance of baseline methods on the forward problem (see Appendix Table 9 for other evaluation metrics). We can see that both Latent U-Net (Large) and Invertible X-Net show better performance than most baselines across all datasets. However, in contrast to the trends observed in the inverse problem, we observe that Invertible X-Net with and without cycle loss shows the best performance compared to Latent U-Net models. We further notice that for simple datasets such as FVA and FFA, FNO has significantly lower MAE compared to other models. We later qualitatively explain this pattern by showing that FNO tends to capture dominant modes of the waveform distribution while neglecting subtler reflection patterns.
@@
Qualitative Comparisons: In Figure 4(a), we compare inverse model predictions on three datasets, CVB, CFB, and STA, choosing one from each family of OpenFWI. We observe that our proposed methods (Latent U-Net and Invertible X-Net) produce considerably better velocity maps than the two SOTA baselines, InversionNet and Auto-Linear. We chose these 3 datasets as they are complex and have high heterogeneity as the velocity rapidly changes in all directions. For heterogeneous velocity maps, the predictions of baseline methods are smooth and they fail to delineate sharp changes in velocities. On the other hand, Latent U-Net and Invertible X-Net are able to predict better velocity fields capturing rapid variations in both shallow and deeper parts (highlighted using black-boxes in the figure). Between Latent U-Net (Large) and Invertible X-Net, we can see that the performance of Latent U-Net is slightly better especially on CFB and STA datasets. Note that Invertible X-Net uses a smaller number of parameters (24M) compared to Latent U-Net (Large) (33M). We provide additional visualizations of more baseline methods and across other datasets in the Appendix.
Figure 4(b) shows visualizations of model predictions for the forward problem on three datasets. Since these datasets are heterogeneous, the recorded seismic waveforms exhibit convoluted interference patterns. In particular, we can see that in all ground truth visualizations, there is a direct arrival in the shallower layers represented as a ‘slanted line’ signature. While capturing the direct arrival is relatively simple, the primary challenge in solving forward problems is to capture the subtler reflections in the deeper layers, where even small magnitudes of errors can result in large differences in inferred velocity maps. We can see that while all baseline methods are able to predict shallow waveforms easily (like direct arrivals), methods such as FNO and Auto-Linear are struggling to predict the seismic waveforms, especially in the deeper regions (highlighted using black-boxes in the Figure). On the other hand, our proposed models consistently improve results over baselines across all datasets. Additional visualizations for the forward problem are provided in the Appendix.
Building on these quantitative and qualitative results, we provide key insights on four open questions in the field of DL4SI in the following.
@@
Figure 5: Effect of latent dimensions and skip connections on the performance of Latent U-Net (Large) (a and b) and Latent U-Net (Small) (c). ‘A’ and ‘B’ show ‘low’ and ‘high’ complexity scenarios for each data group.
@@ 5.1.1 What is the Effect of Latent Space Sizes on Translation Performance?
In Figure 5 (a), we investigate the impact of varying latent space sizes on the quality of domain translations using the Latent U-Net (Large) model on the inverse problem. We compare MAE by reducing latent space dimensions from $70\times 70$ to $8\times 8$. We can see that the quality of translations is not affected much for ‘A’-family of OpenFWI datasets, while the ‘B’-family shows monotonically increasing trend in MAE as we reduce latent dimension. Since the ‘A’-family was intentionally created to represent simpler velocity distributions, even the smallest latent space ($8\times 8$) is sufficient to estimate velocity maps. However, the ‘B’-family of datasets represent more complex velocity distributions, requiring a sufficiently larger latent space for the translation. This indicates that the ideal size of the latent space should be decided based on the complexity of problem being solved.
@@ 5.1.2 Do We Need Complex Architectures for Translations?
We further study the role of model size on tranlslation performance by comparing Latent U-Net (Small) and Latent U-Net (Large) with varying latent dimensions in Figure 5 (b). We find that for a given latent dimension, Latent U-Net (Large) outperforms Latent U-Net (Small), motivating architectures with sufficient complexity for domain translations. We further study the impact of using skip connections in the latent space translation models on varying latent space dimensions in Figure 5 (c).. We can see that adding skip connections shows marginal improvements in performance for larger latent dimensions. However, when the number of dimensions is low (less than 20), we can see a sharp degradation in performance by removing skip-connections. This indicates that smaller latent representations are weaker and thus insufficient for domain translation in U-Net frameworks without being augmented with skip connections.
Figure 6: Comparing a two-stage “Reconstruct then Translate” variant of Latent U-Net (Large) with the default one-stage formulation, across different training fractions.
5.1.3 Should We Train Latent Spaces Solely for Reconstruction?
Auto-Linear and Latent U-Net employ two different schools of thought for learning latent spaces for domain translations. Auto-Linear uses a two-stage “Reconstruct then Translate” approach where both encoder-decoder pairs are first trained solely for reconstruction, followed by the training of latent translation models in the second stage while keeping encoders-decoders frozen. In contrast, Latent U-Net follows a single-stage “Directly Translate” appproach where encoder-decoder pairs are trained together with the translation models, to allow the learned latent spaces to be influenced by the translation objective.
To investigate the effect of using frozen latent spaces on translation performance, we consider a two-stage “Reconstruct then Translate” variant of Latent U-Net where encoder-decoder pairs are first trained for reconstruction followed by translation in Figure 6 for varying OpenFWI families as we reduce the training fraction from 100% to 5%. We can see that for complex families such as CVB and CFB, the MAE of “Reconstruct then Translate” variant is higher for lower training fractions, suggesting that it is useful to influence the training of the latent spaces based on the translation objectives. We see a similar trend for Latent U-Net (Small) shown in Appendix Figure 21.
@@ 5.1.4 Is it Useful to Jointly Solve Forward and Inverse Problems?
As discussed in the quantitative comparisons of Latent U-Net (Large) and Invertible X-Net on the forward problem in Figure 3(b), we see a surprising trend that Invertible X-Net generally outperforms Latent U-Net (Large) despite having smaller number of parameters. To understand this trend, we analyzed the training behavior of Invertible X-Net to learn that X-Net first learns to solve the inverse problem in the initial epochs of training, followed by a gradual shift in learning the solution to the forward problem in later epochs (see Appendix Figures 11 and 12 for visualizations). Since the model optimizes the combined loss on both velocity and waveforms together, as soon as the model is able to achieve a good solution for velocity, the gradients from the combined loss start helping the model to achieve better forward solutions. This also suggests that Invertible X-Net is seemingly learning the connection between forward and inverse problems for effective bidirectional training, and therefore is able to outperform Latent U-Net (Large) despite lower number of model parameters. To verify this, we compared the performance of Latent U-Net (Large), Invertible X-Net and Invertible X-Net trained only in the forward direction (Appendix Figure 10). We observed that indeed Invertible X-Net with forward-only mode shows worst performance followed by Latent U-net (Large), while Invertible X-Net (joint training) shows best performance.
@@ 5.2 Zero-shot Generalization Results
Extrapolating across OpenFWI Datasets: To assess the out-of-distribution generalizability of best-performing models on unseen datasets, in Figure 7, we evaluate zero-shot performance of models trained across all datasets (shown as rows) when tested across all datasets (shown as columns), for both forward and inverse problems. Here, the color intensity indicates difference in SSIM performance of two competing methods. We can see that Invertible X-Net outperforms other baselines for both forward and inverse problems across majority of train-test cases. The only exception is when the models are trained on FVB, a relatively simpler dataset than other data families such as CurveVel and FlatVel. From a geophysics perspective, we expect to have better generalizability for models trained on complex geological settings and tested on relatively simpler ones, as complex geological structures can be thought of as juxtapositions of simpler geological structures. Additionally, we compare the zero-shot generalization for all baselines and Latent U-Net models in detail in the Appendix (Figures 13 - 16).
@@ 6 Conclusions and Future Research Directions
Ethics Statement
@@ Acknowledgments
We are thankful for the support of computational resources provided by the Advanced Research Computing (ARC) Center at Virginia Tech. This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan ( https://www.energy.gov/doe-public-access-plan).
@@ D.3 Model Architecture
We provide more details related to model architecture, layers, and number of parameters related to seismic and velocity encoder-decoder architecture in Table 3, and regarding the latent models used for translation in Table 4. Further, we compare our model parameters with baseline models in Table 6.
@@
Table 6: Comparison of encoder, decoder, and latent model parameters for our model (Latent U-Net and Invertible X-Net) with other baseline models.
@@ E.1 Additional Inverse Modeling Results
Additionally, we also show zero shot generalization of our models on the Marmousi and Overthrust dataset in Tables 8 and 10. Our model Invertible X-Net shows generalizability in SSIM indicating that overall prediction has better geological understanding than other baseline models.
@@
Metric Model FVA FVB FFA FFB CVA CVB CFA CFB STA STB
MAE $\downarrow$ InversionNet 19.67 52.77 102.77 224.57 25.80 158.31 38.90 246.94 93.96 103.37
Auto-Linear 12.16 70.05 110.72 273.02 24.55 181.39 41.38 268.47 107.85 95.63
Latent U-Net 9.01 28.11 48.05 123.56 30.91 92.91 27.05 185.22 67.00 71.41
Invertible X-Net 23.80 34.94 57.86 139.81 59.45 94.11 198.56 181.03 67.09 121.42
MSE $\downarrow$ InversionNet 1002.74 17271.62 36438.33 188044.46 4081.10 68214.92 9490.73 136327.84 23626.76 58622.35
Auto-Linear 1053.03 33907.57 42391.39 236457.48 5952.25 81789.80 13715.42 154713.12 31274.60 21819.14
Latent U-Net 216.92 6464.06 12609.65 91784.50 2348.22 33935.12 3771.38 93249.92 14353.55 14564.94
Invertible X-Net 1245.52 6969.42 14659.46 96121.90 7719.55 32559.95 77105.41 86512.36 13106.96 31550.22
SSIM $\uparrow$ InversionNet 0.9894 0.9461 0.8073 0.6726 0.9765 0.7208 0.9566 0.6136 0.8858 0.6314
Auto-Linear 0.9887 0.9044 0.8056 0.6169 0.97 0.6865 0.9424 0.5695 0.8422 0.7274
Latent U-Net 0.9967 0.9809 0.9273 0.8156 0.991 0.8515 0.98 0.6930 0.9298 0.8064
Invertible X-Net 0.9917 0.9769 0.9135 0.8076 0.9826 0.8532 0.9316 0.7116 0.9360 0.7913

### v3

@@ 5 Results and Discussions
Quantitative Comparisons: Figure 3(a) compares SSIM on velocity predictions while solving the inverse problem across different OpenFWI datasets. We can see that our proposed models, which capture complex non-linear relationships in the latent space using U-Net and IU-Net architectures, consistently outperform baseline methods InversionNet, VelocityGAN, & Auto-Linear across majority of datasets. Overall, the large Latent U-Net model shows the best performance in inverse modeling followed by the small Latent U-Net model. On complex datasets such as Style family and CFB, Invertible X-Net with and without cycle loss are also comparable to the small Latent U-Net model. In addition to SSIM, we also evaluate baselines using other metrics reported in Table 7(Appendix E.1) and compare MAE across models in Appendix figure 3(a).
Figure 3(b) compares the performance of baseline methods on the forward problem (see Appendix E.2 Table 9 for other evaluation metrics). We can see that both Latent U-Net (Large) and Invertible X-Net show better performance than most baselines across all datasets. However, in contrast to the trends observed in the inverse problem, we observe that Invertible X-Net with and without cycle loss shows the best performance compared to Latent U-Net models. We further notice that for simple datasets such as FVA and FFA, FNO has significantly higher SSIM compared to other models. We later qualitatively explain this pattern by showing that FNO tends to capture dominant modes of the waveform distribution while neglecting subtler reflection patterns.
@@
Qualitative Comparisons: In Figure 4(a), we compare inverse model predictions on three datasets, CVB, CFB, and STA, choosing one from each family of OpenFWI. We observe that our proposed methods (Latent U-Net and Invertible X-Net) produce considerably better velocity maps than the two SOTA baselines, InversionNet and Auto-Linear. We chose these 3 datasets as they are complex and have high heterogeneity as the velocity rapidly changes in all directions. For heterogeneous velocity maps, the predictions of baseline methods are smooth and they fail to delineate sharp changes in velocities. On the other hand, Latent U-Net and Invertible X-Net are able to predict better velocity fields capturing rapid variations in both shallow and deeper parts (highlighted using black-boxes in the figure). Between Latent U-Net (Large) and Invertible X-Net, we can see that the performance of Latent U-Net is slightly better especially on CFB and STA datasets. Note that Invertible X-Net uses a smaller number of parameters (24M) compared to Latent U-Net (Large) (33M). We provide additional visualizations of more baseline methods and across other datasets in the Appendix H.
Figure 4(b) shows visualizations of model predictions for the forward problem on three datasets. Since these datasets are heterogeneous, the recorded seismic waveforms exhibit convoluted interference patterns. In particular, we can see that in all ground truth visualizations, there is a direct arrival in the shallower layers represented as a ‘slanted line’ signature. While capturing the direct arrival is relatively simple, the primary challenge in solving forward problems is to capture the subtler reflections in the deeper layers, where even small magnitudes of errors can result in large differences in inferred velocity maps. We can see that while all baseline methods are able to predict shallow waveforms easily (like direct arrivals), methods such as FNO and Auto-Linear are struggling to predict the seismic waveforms, especially in the deeper regions (highlighted using black-boxes in the Figure). On the other hand, our proposed models consistently improve results over baselines across all datasets. Additional visualizations for the forward problem are provided in the Appendix H. Building on these quantitative and qualitative results, we provide key insights on four open questions in the field of DL4SI in the following.
@@
Figure 5: Effect of latent dimensions and skip connections (across latent dimensions) on the performance of Latent U-Net (Large) (a and b) and Invertible X-Net (c). ‘A’ and ‘B’ show ‘low’ and ‘high’ complexity scenarios for each data group.
@@ 5.1.1 What is the Effect of Latent Space Sizes on Translation Performance?
In Figure 5, we investigate the impact of varying latent space sizes on the quality of domain translations using the Latent U-Net (Large) (a) and Invertible X-Net(c) models on the inverse problem. We compare MAE by reducing latent space dimensions from $70\times 70$ to $8\times 8$. We can see that the quality of translations is not affected much for ‘A’-family of OpenFWI datasets, while the ‘B’-family shows monotonically increasing trend in MAE as we reduce latent dimension. Since the ‘A’-family was intentionally created to represent simpler velocity distributions, even the smallest latent space ($8\times 8$) is sufficient to estimate velocity maps. However, the ‘B’-family of datasets represent more complex velocity distributions, requiring a sufficiently larger latent space for the translation. This indicates that the ideal size of the latent space should be decided based on the requirements of the problem being solved.
@@ 5.1.2 Do We Need Complex Architectures for Translations?
We further study the role of model size on translation performance by comparing Latent U-Net (Small) and Latent U-Net (Large) with varying latent dimensions in Figure 18(b) (b). We find that for a given latent dimension, Latent U-Net (Large) outperforms Latent U-Net (Small), motivating architectures with sufficient complexity for domain translations. We further study the impact of using skip connections in the latent space translation models on varying latent space dimensions in Figure 5 (b). We can see that adding skip connections shows marginal improvements in performance for larger latent dimensions. However, when the number of dimensions is low (less than 20), we can see a sharp degradation in performance by removing skip-connections. This indicates that smaller latent representations are weaker and thus insufficient for domain translation in U-Net frameworks without being augmented with skip connections.
Figure 6: Comparison of Latent U-Net (Large) across three training objectives: direct translation, reconstruction followed by translation, and combined learning of both, evaluated at different fractions of training data.
5.1.3 What is the role of manifold learning in DL4SI?
Manifold learning for latent spaces in domain translation can follow three strategies: (1) the “Reconstruct then Translate” (Auto-Linear), where encoder-decoder pairs are first trained for reconstruction, followed by translation on frozen latent spaces; (2)“Directly Translate” strategy (Latent U-Net), which jointly trains encoder-decoder pairs and translation models to optimize latent spaces for translation; and (3) a hybrid “Reconstruct and Translate” approach, where the model is trained with both reconstruction and translation losses. To investigate their impact, we evaluate these approaches across varying training data fractions for OpenFWI families (Figure 6). Our findings reveal that the “Directly Translate” approach consistently outperforms the other two, highlighting that latent spaces optimized solely for translation are most effective. Hence, manifold learning, while beneficial for reconstruction, may lead to suboptimal performance when translation is the primary goal. We see a similar trend for Latent U-Net (Small) and Invertible X-Net shown in Appendix F.3 Figure 22.
@@ 5.1.4 Is it Useful to Jointly Solve Forward and Inverse Problems?
As discussed in the quantitative comparisons of Latent U-Net (Large) and Invertible X-Net on the forward problem in Figure 3(b), we see a surprising trend that Invertible X-Net generally outperforms Latent U-Net (Large) despite having smaller number of parameters. Analysis of its training behavior reveals that Invertible X-Net initially focuses on solving the inverse problem, gradually shifting to the forward problem in later epochs. (see Appendix E.3 Figures 12 and 13 for visualizations). This is facilitated by the combined loss function, where gradients from the inverse problem enhance forward problem optimization once a good velocity solution is achieved. This bidirectional learning enables Invertible X-Net to outperform Latent U-Net (Large). Further comparisons in Appendix E.3 confirm that Invertible X-Net trained only on the forward problem performs worse than both Latent U-Net (Large) and the jointly trained Invertible X-Net, emphasizing the advantage of joint optimization.With average ranks of 2.7 (MAE) and 1.9 (SSIM) for inverse problems, and 1.4 (MAE) and 1.1 (SSIM) for forward problems (Appendix Tables 7 and 9), Invertible X-Net offers a balanced and efficient solution for joint tasks with just 26M parameters. In contrast, Latent U-Net, while slightly better for inverse-only tasks, requires 70M parameters for both forward and inverse problems (35M x2), making Invertible X-Net the more practical and effective choice for the combined optimization.
@@ 5.2 Zero-shot Generalization Results
Extrapolating across OpenFWI Datasets: To assess the out-of-distribution generalizability of best-performing models on unseen datasets, in Figure 7, we evaluate zero-shot performance of models trained across all datasets (shown as rows) when tested across all datasets (shown as columns), for both forward and inverse problems. Here, the color intensity indicates difference in SSIM performance of two competing methods. We can see that Invertible X-Net outperforms other baselines for both forward and inverse problems across majority of train-test cases. The only exception is when the models are trained on FVB, a relatively simpler dataset than other data families such as CurveVel and FlatVel. From a geophysics perspective, we expect to have better generalizability for models trained on complex geological settings and tested on relatively simpler ones, as complex geological structures can be thought of as juxtapositions of simpler geological structures. Additionally, we compare the zero-shot generalization for all methods in detail in Appendix E.4 (Figures 14 - 17).
@@ 6 Conclusions and Future Research Directions
7 Reproducibility Statement
All the code required to train and evaluate the proposed methods, as well as the baselines, has been uploaded to an anonymous GitHub repository: https://github.com/KGML-lab/Generalized-Forward-Inverse-Framework-for-DL4SI. The data and corresponding processing code used in this work are sourced from the OpenFWI website: https://openfwi-lanl.github.io/docs/data.html. The dataset is further described in Appendix Section B and Table 2. The model architecture and hyper-parameter details for all proposed models and baselines are thoroughly discussed in Appendix Section D.
8 Ethics Statement
@@ Acknowledgments
This work was supported in part by NSF awards IIS-2239328 and IIS-2107332. We are grateful to the Advanced Research Computing (ARC) Center at Virginia Tech for providing access to GPU compute resources for this project. This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan ( https://www.energy.gov/doe-public-access-plan).
@@ D.3 Model Architecture
We summarize details related to model architecture, layers, and number of parameters related to seismic and velocity encoder-decoder architecture in Table 3, and regarding the latent models used for translation in Table 4. Further, we compare our model parameters with baseline models in Table 6.
D.3.1 Latent U-Net
Latent U-Net architectures are specifically adapted to subsurface imaging by employing U-Net-based encoder-decoder pairs designed to handle domain-specific input and output dimensions, such as velocity maps and seismic waveforms, while enabling latent dimension processing. Both the Large and Small variants utilize identical encoder and decoder designs, featuring 5 layers in the encoder and 6 layers in the decoder. These layers operate on an embedding dimension of 128x70x70, with channel sizes ranging from 8 to 128 in the encoder and 128 to 1 (velocity) or 128 to 5 (waveform) in the decoder.
The Large Latent U-Net incorporates 2 depth levels and 4 convolutional blocks per depth level in its latent space translation model, resulting in a total parameter size of 34.96M. The Small variant, designed for reduced complexity, includes 1 convolutional block per depth level in its latent translation model, reducing the total parameter size to 18.13M.
D.3.2 Invertible X-Net
Invertible X-Net leverages an iUNet-based architecture to achieve bidirectional mappings, enabling consistent forward (velocity-to-waveform) and inverse (waveform-to-velocity) transformations within a shared framework. iUNet introduces invertible coupling blocks, ensuring bijectivity for the latent space translation model. Max pooling is replaced with orthogonal convolutional filters for downsampling, while upsampling is performed using orthogonal deconvolution filters to maintain invertibility.
While the latent space translation model in Invertible X-Net is fully invertible due to its iUNet structure, the encoder-decoder pairs are not invertible. Therefore, Invertible X-Net as a whole is not fully invertible. However, by employing two separate encoder-decoder pairs for velocity and waveform domains, Invertible X-Net achieves an architecture that enables bidirectional mappings between the two domains while preserving consistency.
Invertible X-Net features 4 depth levels, with 4 invertible coupling blocks per level, and adopts an encoder-decoder structure similar to Latent U-Net, with 5 encoder layers and 6 decoder layers, an embedding dimension of 128x70x70, and channel sizes ranging from 8 to 128. The latent translation model - iUNet consists of 25.78M parameters, contributing to a total model size of 26.06M.
@@ D.3.2 Invertible X-Net
Table 6: Comparison of encoder, decoder, and latent model parameters for our model (Latent U-Net and Invertible X-Net) with other baseline models. The parameters for Latent U-Nets and Invertible X-Nets are calculated for Latent dimension 70.
@@ E.1 Additional Inverse Modeling Results
Additionally, we also show zero shot generalization of our models on the Marmousi and Overthrust dataset in Tables 8. Our model Invertible X-Net shows generalizability in SSIM indicating that overall prediction has better geological understanding than other baseline models.
@@
Metric Model FVA FVB CVA CVB FFA FFB CFA CFB STA STB Average Rank
MAE $\downarrow$ InversionNet 19.67 52.77 102.77 224.57 25.80 158.31 38.90 246.94 93.96 103.37 2.8
Auto-Linear 12.16 70.05 110.72 273.02 24.55 181.39 41.38 268.47 107.85 95.63 3.5
Latent U-Net 9.01 28.11 48.05 123.56 30.91 92.91 27.05 185.22 67.00 71.41 1.3
Invertible X-Net 23.80 34.94 57.86 139.81 59.45 94.11 198.56 181.03 67.09 121.42 2.7
MSE $\downarrow$ InversionNet 1002.74 17271.62 36438.33 188044.46 4081.10 68214.92 9490.73 136327.84 23626.76 58622.35 2.8
Auto-Linear 1053.03 33907.57 42391.39 236457.48 5952.25 81789.80 13715.42 154713.12 31274.60 21819.14 3.5
Latent U-Net 216.92 6464.06 12609.65 91784.50 2348.22 33935.12 3771.38 93249.92 14353.55 14564.94 1.3
Invertible X-Net 1245.52 6969.42 14659.46 96121.90 7719.55 32559.95 77105.41 86512.36 13106.96 31550.22 2.4
SSIM $\uparrow$ InversionNet 0.9894 0.9461 0.8073 0.6726 0.9765 0.7208 0.9566 0.6136 0.8858 0.6314 3
Auto-Linear 0.9887 0.9044 0.8056 0.6169 0.97 0.6865 0.9424 0.5695 0.8422 0.7274 3.8
Latent U-Net 0.9967 0.9809 0.9273 0.8156 0.991 0.8515 0.98 0.6930 0.9298 0.8064 1.3
Invertible X-Net 0.9917 0.9769 0.9135 0.8076 0.9826 0.8532 0.9316 0.7116 0.9360 0.7913 1.9